OpenStudy (anonymous):

Which reason supports the true statement that angle1 and angle8 are supplementary? A. if angle1 and angle8 are same-side interior angles. B. angle1 and angle8 are corresponding angles. C. angle1 and angle8 are a linear pair. D. angle1 and angle8 are vertical angles.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

What do you think the answer is?

OpenStudy (anonymous):

well i got C

OpenStudy (ranga):

Yes!

OpenStudy (anonymous):

.If angle9 = 7y° and angle10 = 3y°, what is the measure of angle16 such that x is parallel to z? A. 54° B. 90° C. 126° D. 18°

OpenStudy (anonymous):

this one i got D

OpenStudy (ranga):

You solved for y which is 18 but that is not what they are asking.

OpenStudy (ranga):

Having solved for y how much is angle 9?

OpenStudy (anonymous):

it says 9=7.

OpenStudy (ranga):

angle 9 is 7 times y.

OpenStudy (ranga):

Let us start from the beginning. What is the sum of angles 9 and 10?

OpenStudy (anonymous):

isnt it 7 and 3? or am i suppose to look for something else.

OpenStudy (ranga):

The problem says angle 9 = 7y and angle 10 = 3y So what is the sum of angles 9 and 10?

OpenStudy (ranga):

y is an unknown that we have to solve for.

OpenStudy (anonymous):

how do i find the sum?

OpenStudy (ranga):

Add 7y and 3y. 7y + 3y = ?

OpenStudy (ranga):

just simplify the expression 7y + 3y

OpenStudy (ranga):

Example: 5x + 3x = 8x

OpenStudy (anonymous):

10

OpenStudy (ranga):

It is 10y

OpenStudy (ranga):

Sum of angle 9 and angle 10 = 10y Looking at the figure do you recognize angles 9 and 10 are supplementary?

OpenStudy (anonymous):

yeah

OpenStudy (ranga):

And supplementary angles always add up to how many degrees?

OpenStudy (anonymous):

180

OpenStudy (ranga):

Yes. Previously we found the sum of angle 9 and angle 10 = 10y. And because angle 9 and angle 10 are supplementary they add up to 180 So 10y = 180 What is y?

OpenStudy (ranga):

Divide both sides by 10

OpenStudy (anonymous):

18

OpenStudy (ranga):

yes y = 18. angle 9 = 7 x y = 7 x 18 = ?

OpenStudy (anonymous):

is that 7 times y or is it 7x?

OpenStudy (ranga):

seven times

OpenStudy (ranga):

angle 9 = (7)(y) = (7)(18) = ?

OpenStudy (anonymous):

im not really sure

OpenStudy (ranga):

you just need to multiply 7 and 18

OpenStudy (anonymous):

oooh 126

OpenStudy (ranga):

Yes. If angle 9 is 126 what will angle 14 be?

OpenStudy (ranga):

Examples: In the diagram, angles 1 and 6 are called vertical opposite angles and they are equal. Angles 2 and 5 are vertical opposites and they are equal. Angles 3 & 8 are equal. Angles 4 & 7 are equal. Now can you answer the previous question?